One month ago you bought some shares of stock at $123 per share. Today the stock is worth $130 per share. What is the continuously-compounded monthly return for this period?

Enter answer in percents, accurate to two decimal places.
Answer #1

here as we know that future value =present value*exp(rt) where t is number of periods and r is rate of return

hence 130=123*exp(r*1)

r =ln(130/123)=0.05535 ~ 5.54%